OpenStudy (anonymous):

my sister is 14 years older than I am. the sum of our ages is 30. find our ages

OpenStudy (anonymous):

she's 26, ur 14

OpenStudy (anonymous):

maybe ? idk . i just confused myself

OpenStudy (anonymous):

x+y=30

OpenStudy (anonymous):

y = x+30

OpenStudy (anonymous):

solve these two

OpenStudy (anonymous):

i am sorry

OpenStudy (anonymous):

the second eqn is y= x+14

OpenStudy (lalaly):

let x be ur age x+14 is ur sisters age sum of ur ages x+x+14 = 30 2x=30-14 2x=16 x=8 your 8 shes 22

OpenStudy (anonymous):

30-14=16 so one of u is 16 and the other is x